# GeoLRM: Geometry-Aware Large Reconstruction Model for High-Quality 3D Gaussian Generation

### v1

@@ 4.5 Ablation Study
We provide ablation studies for the key designs of our method as shown in Table 3. Due to the limited computation sources, the ablation is done using a smaller reconstruction model (12 layers) and lower resolution (224x224).
Table 3: Ablation study. Models are tested on the GSO dataset [13]. Upper: 6 input views and 4 testing views. Lower: different input views. Bold and underline denote the highest and second-highest scores, respectively.
@@ 4 Inputs 8 Inputs 12 Inputs
Mothod PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
@@
Hierarchical Image Encoder Our ablation study underscores the critical role of hierarchical image features in reconstruction tasks, which necessitate both high-level semantic information (e.g., object identity and arrangement) and low-level texture information (e.g., surface patterns and colors). As illustrated in Figure 5, the absence of high-level features leads to model instability, while omitting low-level features results in a loss of textural detail. This dual requirement emphasizes the model’s reliance on a comprehensive feature set for accurate image reconstruction.
3D RoPE In transformer-based architectures, the role of positional embeddings is critical for accurately interpreting sequence data positions. A key question arises: With the reconstruction transformer employing deformable cross attention to elevate 2D features to 3D, is positional embedding still necessary? Our ablation studies confirm its necessity. Notably, 3D RoPE significantly enhances the model’s ability to handle longer sequences. For instance, increasing the sequence length from 4k to 16k elements, models equipped with 3D RoPE exhibited a PSNR improvement of 0.4, compared to a 0.2 improvement in models lacking 3D RoPE. This observation aligns with the 1D RoPE [52].
Dynamic Input The ablation study demonstrates a decrease in performance when employing our dynamic input view strategy compared to the fixed 6 input view setting when the training and testing phases were consistent. Despite this, the dynamic input strategy enhances the model’s ability to generalize across different input configurations. This adaptability is critical for handling more complex scenarios, aligning with our primary objectives.
5 Limitations
Although our two-stage method achieves high-quality reconstruction, it is not an end-to-end model, which leads to error accumulation. In other words, the results of the second stage highly depend on the occupancy grids of the first stage. Currently, we have to use the proposal network since directly processing Gaussian points in the whole 3D space is time-consuming. For the future work, we plan to extend our method in an end-to-end manner.
@@ 6 Conclusion
In this paper, we propose geometry-aware large reconstruction model for efficient and high-quality 3D generation. Different with previous works, our method explores the sparsity of 3D and leverages the explicit geometric relationship between 3D and 2D images. To this end, GeoLRM adopts a 3D-aware transformer structure to predict 3D Gaussians in a coarse-to-fine fashion. Specifically, we first utilize a proposal network to predict the coarse occupancy grids of 3D assets, which provide the initial 3D anchor points for the second stage. Then we employee the deformable cross attention to refine the 3D structure. Experimental results show that the proposed method can efficiently process higher resolution and denser image inputs with better performance.
@@ Appendix A Appendix
A.1 Mesh Extraction from 3D Gaussians
We follow the mesh extraction pipeline from [54] to extract high quality mesh representations from 3D Gaussians. The result is shown in Figure 6.
Figure 6: Image-to-3D generation with mesh extraction results.
A.2 Occupancy Ground Truth
@@ A.3 More Implementation Details
We illustrate the details of network architecture and training procedure in Table 4. We train both the proposal transformer and the reconstruction transformer for 12 epochs on GObjaverse [41], which takes 0.5 and 2 days respectively on 32 A100 40G. For the proposal transformer, we use a batch size of 2 per GPU and apply mixed-precision training with BF16 data type. For the reconstruction transformer, we use a batch size of 1 per GPU and keep the full precision. We note that the second stage is particularly sensitive to the data type and would fail if using mixed-precision.
Table 4: Implementation details.

### v2

@@ 4.5 Ablation Study
In this part, We provide ablation studies for the key designs of our method as shown in Table 4. Due to the limited computational sources, the ablation is done using a smaller reconstruction model (12 layers) and lower resolution (224x224).
Table 4: Ablation study of some key designs. Models are tested on the GSO dataset [13]. Upper: 6 input views and 4 testing views. Lower: different input views. Bold and underline denote the highest and second-highest scores, respectively.
@@ 4 Inputs 8 Inputs 12 Inputs
Method PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
@@
Hierarchical Image Encoder Our ablation study underscores the critical role of hierarchical image features in reconstruction tasks, which necessitate both high-level semantic information (e.g., object identity and arrangement) and low-level texture information (e.g., surface patterns and colors). As illustrated in Figure 5, the absence of high-level features leads to model instability, while omitting low-level features results in a loss of textural detail. This dual requirement emphasizes the model’s reliance on a comprehensive feature set for accurate image reconstruction. We also performed an ablation study regarding the Plücker ray embeddings in the low-level encoder. These coordinates assist the model in learning camera directions, contributing to an improved performance.
3D RoPE In transformer-based architectures, the role of positional embeddings is critical for accurately interpreting sequence data positions. A key question arises: With the reconstruction transformer employing deformable cross-attention to elevate 2D features to 3D, is positional embedding still necessary? Our ablation studies confirm its necessity. Notably, 3D RoPE significantly enhances the model’s ability to handle longer sequences. For instance, increasing the sequence length from 4k to 16k elements, models equipped with 3D RoPE exhibited a PSNR improvement of 0.4, compared to a 0.2 improvement in models lacking 3D RoPE. This observation aligns with the 1D RoPE [52].
Dynamic Input The ablation study demonstrates a decrease in performance when employing our dynamic input view strategy compared to the fixed 6-input view setting when the training and testing phases were consistent. Despite this, the dynamic input strategy enhances the model’s ability to generalize across different input configurations. This adaptability is critical for handling more complex scenarios, aligning with our primary objectives.
Deformable attention As shown in Table 5, the ablation results indicate that increasing the number of sampling points in the deformable attention generally improves performance. Given the trade-off between computational cost and performance gain, we find that using 8 sampling points strikes the best balance.
Table 5: Ablation study of deformable attention. ‘0 sampling points’ means directly using the projected points without any deformation. Bold and underline denote the highest and second-highest scores, respectively.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
0 sampling points 19.52 0.802 0.265
4 sampling points 20.21 0.819 0.238
8 sampling points 20.73 0.839 0.220
16 sampling points 20.80 0.846 0.219
@@ 5 Conclusion
In this paper, we present GeoLRM, a geometry-aware large reconstruction model designed to improve the efficiency and quality of 3D generation. Our approach distinguishes itself from previous methods by effectively utilizing the inherent sparsity of 3D structures and explicitly integrating geometric relationships between 3D and 2D images. The GeoLRM framework employs a 3D-aware transformer architecture that predicts 3D Gaussians through a sophisticated coarse-to-fine methodology. Initially, a proposal network estimates coarse occupancy grids, which serve as foundational 3D anchor points for subsequent refinement. The second stage leverages deformable cross-attention to enhance the 3D structure, integrating detailed textural information. Extensive experiments validate that GeoLRM can process higher resolutions and accommodate denser image inputs, outperforming existing models in terms of detail and accuracy. This innovation demonstrates significant potential for real-world applications, particularly in domains where dense view inputs can enhance output quality and user experience. GeoLRM’s ability to handle up to 21 images efficiently underscores its scalability and adaptability, paving the way for integration with advanced video generation technologies.
6 Limitation
While GeoLRM achieves impressive reconstruction quality, it does so through a two-stage process, which is not inherently end-to-end. This segmentation can lead to the accumulation of errors. The reliance on a proposal network is currently indispensable due to the computational intensity of processing Gaussian points across the entire 3D space. This necessity introduces potential inefficiencies and constraints that could hinder real-time applications. Future research will focus on developing an end-to-end solution that integrates these stages seamlessly, reducing error propagation and optimizing processing time. By addressing these limitations, we aim to enhance the model’s robustness and applicability across a broader range of 3D generation tasks.
Acknowledgement
This work was supported in part by the Beijing Natural Science Foundation under Grant No. L247009 and in part by Young Elite Scientists Sponsorship Program by CAST (No. 2024QNRC003).
@@ Appendix A Appendix
A.1 Occupancy Ground Truth
@@
Figure A: Image-to-3D generation with mesh extraction results.
A.2 Mesh Extraction from 3D Gaussians
We adopt the mesh extraction pipeline from [54] to derive high-quality mesh representations from 3D Gaussians. Figure A illustrates the mesh generation results of our method, while Figure B compares our generated mesh with other techniques. The results demonstrate the effectiveness of our approach, despite some loss of detail during conversion.
Figure B: Comparison of the generated meshes.
@@ A.3 More Implementation Details
We illustrate the details of network architecture and training procedure in Table A. We train both the proposal transformer and the reconstruction transformer for 12 epochs on GObjaverse [41], which takes 0.5 and 2 days respectively on 32 A100 40G. For the proposal transformer, we use a batch size of 2 per GPU and apply mixed-precision training with BF16 data type. For the reconstruction transformer, we use a batch size of 1 per GPU and keep the full precision. We note that the second stage is particularly sensitive to the data type and would fail if using mixed-precision.
Table A: Implementation details.
@@
A.4 Social Impact
3D AIGC is transforming sectors by automating realistic 3D model creation. In entertainment, it streamlines film and game production, reducing costs and enhancing experiences. Education benefits from immersive VR simulations for deeper learning. Architecture sees rapid design visualization and urban planning improvements. Challenges include job displacement and ethical concerns over content authenticity. Addressing these requires legal and policy measures, such as clear copyright laws and standards to protect intellectual property. Developing advanced content moderation tools can detect false content, and enhancing AI security can prevent misuse. By focusing on these solutions, we can mitigate negative impacts and maximize the positive contributions of 3D AIGC to society.